OpenStudy (anonymous):

Does anyone know the rates of change of those two equations? 2x + 3y = −9 y = 3x − 6

OpenStudy (anonymous):

@TheSmartOne @wio please help?

OpenStudy (anonymous):

it means the slope

OpenStudy (anonymous):

the slope of \[y=mx+b\] is \(m\) so the slope of \(y=3x-6\) is \(3\)

OpenStudy (anonymous):

the slope of \[ax+by=c\] is \[m=-\frac{a}{b}\]

OpenStudy (anonymous):

@satellite73 wait is the slope of 2x + 3y = -9 is it positive or negative

OpenStudy (anonymous):

@Nnesha 2x + 3y = -9 is the slope positive or negative

Nnesha (nnesha):

slope intercept form is y = mx +b here m is slope so you have to sole for y to do that you have to subtract 2x both side when you subtract 2x both side what you should get ??

OpenStudy (anonymous):

3x = 11?

Nnesha (nnesha):

nope you cannot combine -2x and -9 bec they are not like terms -2 have x and -9 doesn't have any variable so just write  now divide both side by 3